# Safe MPC Alignment with Human Directional Feedback

### v1

@@ IV-C Properties and Geometric Interpretations
In this session, we present some properties and the geometric interpretations for the proposed safe MPC alignment method. The properties will be used to establish the convergence of the algorithm in the next section.
@@ IV-C1 Piecewise Linearity of Cuts
Recall the weight-feature parameterization of the safety constraint (5). From (12), we can examine $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$ is a linear inequality in $\boldsymbol{\theta}$.
Based on the definition of the penalty objective function (2), $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is a linear fractional inequality
with a positive denominator $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ due to (12). Multiplying the denominator on both side yields another linear inequality. Formally, we give a lemma here about the piecewise linearity of the cut $\mathcal{C}_{i}$ in (12).
@@
Equation (15) can be directly obtained from (12), (2) and (5). This completes the proof. ∎
With the lemma above, the intersection $\Theta_{i-1}\cap\mathcal{C}_{i}$ can be viewed as cutting $\Theta_{i-1}$ with the boundary of $\mathcal{C}_{i}$ which consists of two piecewise hyperplanes:
@@
which is shown in Fig. 4
As the initial hypothesis space $\Theta_{0}\in\mathbb{R}^{r}$ in (10) is a polytope. Iteratively cutting with piecewise hyperplanes yields a series of convex polytopes in $\mathbb{R}^{r}$. Consequently, all $\Theta_{i}$s, $i=0,1,2,...$, are convex polytopes.
@@ IV-C2 Relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$
According to (15), the cut of $\mathcal{C}_{i}$ depends on robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ solved by robot MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ in (3) with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. This suggests some relationship between $\boldsymbol{\theta}_{i}$ and the boundary $\mathcal{C}_{i}$ in (18). In fact, the relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$ is given by following lemma.
@@
$\forall i,\ \boldsymbol{\theta}_{i}$ is located on $\mathbf{bd}\ \mathcal{C}_{i}$ in (18). Specifically, $\boldsymbol{\theta}_{i}$ satisfies
@@
See Appendix A-B. ∎
The above lemma states $\boldsymbol{\theta}_{i}\in\mathbf{bd}\ \mathcal{C}_{i}$. Its geometric interpretation is in Fig. 5: the piecewise cut of $\mathcal{C}_{i}$ always passes $\boldsymbol{\theta}_{i}$. This suggests that the choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ is critical in the learning process: it directly affects the position of the cut $\mathbf{bd}\ \mathcal{C}_{i}$ and consequently the volume to be cut out. Thus, it is important to choose a good $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ such that a large hypothesis space is removed after each cut.
We here briefly discuss the implication of different choices of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. When choosing $\boldsymbol{\theta}_{i}$ near the boundary of $\Theta_{i-1}$, as shown in Fig. 5 and Fig. 5, the removed volume of the hypothesis space depends on the specific direction of cut $\mathbf{bd}\,\mathcal{C}_{i}$. Since the direction of the cut $\mathbf{bd}\,\mathcal{C}_{i}$ will be eventually dependent upon the user’s correction direction according to (18), the removed volume of the hypothesis space can be relatively small, as in Fig. 5, or large, as in Fig. 5, depending on the specific direction of the correction applied by the user.
As the specific direction of user correction is hard to assume, one can intuitively see that one good choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ will be near the center of $\Theta_{i-1}$. As such, we can ensure an averaged large cut of the hypothesis space regardless of the direction of $\mathbf{bd}\ \mathcal{C}_{i}$, as shown in Fig. 5. We will rigorously discuss the centering choice of $\boldsymbol{\theta}_{i}$ and the guarantee for volume reduction rate of the hypothesis space in the next section.
Figure 5: Illustration of the different choices of $\boldsymbol{\theta}_{i}$. (a) and (b): When $\boldsymbol{\theta}_{i}$ is chosen near the set boundary, the cut volume (i.e., the volume of the removed hypothesis space) depends on the specific direction of $\mathbf{bd}\ \mathcal{C}_{i}$ (i.e., the user-specific directional correction). (c): When $\boldsymbol{\theta}_{i}$ is near the center, the cut volume is large on average regardless of the specific user directional correction.
V Implementation Details and Main Results
In this section, we will present the details for how to select $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ at each iteration $i$ of the hypothesis space cutting in our proposed Safe MPC Alignment algorithm. Based on the preceding geometric intuition, our choice of $\boldsymbol{\theta}_{i}$ is to achieve a large averaged reduction rate for hypothesis space volume given human directional corrections. The reduction rate enables us to establish an exponential convergence of the learning process. As a result, we can find an upper bound of total user corrections to guarantee the success of learning, i.e., $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$, or identify the misspecification of hypothesis space.
@@ V-A The Choice of $\boldsymbol{\theta}_{i}$
As discussed previously, $\boldsymbol{\theta}_{i}$ needs to be picked from the "center" of $\Theta_{i-1}$ in the proposed algorithm. An apparent option is to choose the mass center of $\Theta_{i-1}$. However, computing the centroid of a polytope can be #P-Hard [56]. The methods, such as Hit-And-Run algorithm [57], which estimate the mass center by sampling, also suffer from the high sample complxity, thus are impractical in our online real-time setting. In [58], the authors provide an another method to choose the center of a polytope using a set of inscribed balls, but it lacks of the guarantee of the volume reduction rate. For computational efficiency and guaranteed convergence, we will use the center of the Maximum Volume Ellipsoid (MVE) inscribed in the convex set $\Theta_{i-1}$ as $\boldsymbol{\theta}_{i}$, shown in Fig. 6.
@@
=$\mathcal{E}=\{\bar{H}\boldsymbol{\theta}+\bar{\boldsymbol{d}}\ |\ \|\boldsymbol{\theta}\|_{2}\leq 1\}.$ (20)
The affine transformation parameter $\bar{H}\in\mathbb{S}^{r}_{++}$ (a $r\times r$ positive definite matrix), and $\bar{\boldsymbol{d}}\in\mathbb{R}^{r}$ is called the center of $\mathcal{E}$. $\bar{H},\bar{\boldsymbol{d}}$ are determined via the following optimization problem for the volume of the ellipsoid $\det H$:
$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\max}\quad&\log\det H\\ \rm{s.t.}\quad&\sup\nolimits_{\|\boldsymbol{\theta}\|_{2}\leq 1}\mathbbm{1}_{\Theta}(H\boldsymbol{\theta}+\boldsymbol{d})\leq 0\end{split},$ (21)
where the indicator function $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=0$ for $\boldsymbol{\theta}\in\Theta$ and otherwise $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=+\infty$. Specifically, when $\mathbbm{1}_{\Theta}$ is a polytope $\{\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i},\ i=1,\dots,m\}$, the problem can be converted to a convex optimization problem with $m$ constraints:
$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\min}\quad&\log\det H^{-1}\\ \rm{s.t.}\quad&\left\lVert H\boldsymbol{h}_{i}\right\rVert_{2}+\boldsymbol{d}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i},\ i=1,\dots,m\end{split}.$ (22)
In the Safe MPC Alignment algorithm, by the above definition, let $\mathcal{E}_{i-1}$ denote the MVE of $\Theta_{i-1}$ and $\bar{\boldsymbol{d}}_{i-1}$ denote the center of $\mathcal{E}_{i-1}$. Then, the choice of $\boldsymbol{\theta}_{i}$ is
@@
Given $\Theta_{i-1}$ a convex polytope, the optimization problem (21) can be solved efficiently by convex programming (22).
@@ V-B Result 1: Finite Iteration Convergence
In this section, we give the main result that, when using the choice of $\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}$ mentioned above, the Safe MPC Alignment algorithm can converge in a finite number of iterations to find $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subset\Theta_{H}$. This result comes from the fact that the volume reduction rate after each hypothesis space cut is upper bounded, as formally stated below.
@@
In the Safe MPC Alignment algorithm, with $\boldsymbol{\theta}_{i}$ as the center $\boldsymbol{\bar{d}}_{i-1}$ of MVE of $\Theta_{i-1}$, one has
$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r}.$ (24)
@@
See Appendix A-C. ∎
The lemma above asserts an upper bound of the volume of the updated hypothesis space $\mathbf{Vol}(\Theta_{i})$ relative to the previous volume. With the above result, we can now establish the finite iteration convergence of the proposed Safe MPC Alignment, stated below.
@@
In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center $\boldsymbol{\bar{d}}_{i-1}$ of MVE of $\Theta_{i-1}$, $i=1,2,3,...$. If $\bar{\Theta}_{H}\subset\Theta_{0}$, the algorithm will terminate successfully within
@@
See Appendix A-D. ∎
The above main result asserts a maximum iteration count for the proposed algorithm to find $\boldsymbol{\theta}_{i}\in\Theta_{H}$. In other words, within $K$ rounds of human corrections, the robot can successfully learn a safety set intended by the user. The above result requires a pre-set of $\rho_{H}^{r}$, which is the radius of a ball contained in $\bar{\Theta}_{H}$. Since the size of $\bar{\Theta}_{H}$ is unknown in practice, $\rho_{H}>0$ can be interpreted as an underestimated size of the unknown user intent set $\bar{\Theta}_{H}$. It also serves as a termination threshold for learning convergence: a smaller $\rho_{H}$ can lead to a more concentrated learned $\boldsymbol{\theta}$.
However, the finite iteration convergence certificate in Theorem 1 comes with a prerequisite: $\bar{\Theta}_{H}\in\Theta_{0}$ has to be satisfied in the first place in the Safe MPC Alignment algorithm. Although such a condition can be achieved by empirically choosing $\Theta_{0}$ large enough to contain all possible $\bar{\Theta}_{H}$ (10), it is also desired to have a systematic way to detect the hypothesis misspecification. In the next session, we will discuss the identification of hypothesis misspecification, where such condition is not met.
@@ V-C Result 2: Hypothesis Misspecification Declaration
We consider the case of hypothesis space misspecification, where $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$. In this case, no $\boldsymbol{\theta}\in\Theta_{0}$ can account for the user intended safety (4). As mentioned in [54, 60], this kind of gap between hypothesis space and true human model can lead to undesired behavior of the robot. The following result states the proposed Safe MPC Alignment algorithm can automatically certify the misspecification of the initial hypothesis space.
Theorem 2 (Identifying misspecification).
In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center $\boldsymbol{\bar{d}}_{i-1}$ of MVE of $\Theta_{i-1}$, $i=1,2,3,...$. If $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the algorithm converges to a set $\Theta^{*}$ as $i\rightarrow\infty$, such that $\mathbf{Vol}(\Theta^{*})=0$ and $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$.
@@
See Appendix A-E. ∎
The above result states that, when the initial hypothesis space is misspecified, $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the Safe MPC Alignment algorithm will finally converge to a set of zero volume on the boundary of the initial hypothesis space $\Theta_{0}$.
The above theorem inspires us to empirically identify the misspecification by examining the position of MVE centers during the converging process: if after the maximum iteration step $K-1$ in (25), the MVE center $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of initial hypothesis set (10), there is likely hypothesis space misspecification and further adjustment on initial hypothesis space should be made. See Fig. 7.
Figure 7: When there is misspecification, the Safe MPC Alignment algorithm will converge to some point on the boundary of $\Theta_{0}$. So we can use the distance from $\boldsymbol{\theta}_{K-1}$ to the boundary as a signal for misspecification.
Specifically, recall (10) that $\Theta_{0}$ is a box with bounding vector $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$, we can pre-define a small threshold $\epsilon$ to examine every entry of $\boldsymbol{\theta}_{K-1}$ and identify possible misspecification. If $\boldsymbol{\theta}_{K-1}$ satisfies the criterion
$\min_{j}\ \lvert\boldsymbol{c}_{l}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert\leq\epsilon\ \text{or}\ \min_{j}\ \lvert\boldsymbol{c}_{h}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert\leq\epsilon,$ (26)

### v2

@@ IV-C Properties and Geometric Interpretations
In this session, we present some properties and the geometric interpretations for the proposed safe MPC alignment. Those properties will be used to establish the convergence of the algorithm in the next section.
@@ IV-C1 Piecewise Linearity of Cuts
By recalling the weight-feature parameterization of the safety constraint (5), one can examine that the inequities in (12) are both linear in $\theta$. This means that the cutting set $\mathcal{C}_{i}$ in (12) is a piecewise linear cut, as stated formally in the following lemma.
@@
See Appendix A-B. ∎
According to the above lemma, $\Theta_{i-1}\cap\mathcal{C}_{i}$ can be viewed as cutting $\Theta_{i-1}$ with the boundary of $\mathcal{C}_{i}$ which is two hyperplanes:
@@
shown in Fig. 4.
As the initial hypothesis space $\Theta_{0}\in\mathbb{R}^{r}$ in (10) is a convex polytope, iteratively cutting with hyperplane pairs yields a series of convex polytopes in $\mathbb{R}^{r}$. Consequently, all $\Theta_{i}$s, $i=0,1,2,...$, are convex polytopes.
@@ IV-C2 Relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$
According to (14), the cutting set $\mathcal{C}_{i}$ depends on plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ which is a solution to penalty-based MPC (3) with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. The following assertion states a relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$.
@@
$\forall i,\ \boldsymbol{\theta}_{i}$ is on $\mathbf{bd}\ \mathcal{C}_{i}$ in (17). Specifically, $\boldsymbol{\theta}_{i}$ satisfies
@@
See Appendix A-C. ∎
Lemma 4 states $\boldsymbol{\theta}_{i}\in\mathbf{bd}\ \mathcal{C}_{i}$. Its geometric interpretation is in Fig. 5: the cut $\mathbf{bd}\ \mathcal{C}_{i}$ always passes $\boldsymbol{\theta}_{i}$. This suggests that the choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ is critical in the learning process: it directly affects the position of the cut $\mathbf{bd}\ \mathcal{C}_{i}$ and consequently the removal volume. It is important to choose a good $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ such that a large hypothesis space is removed after each cut.
We briefly discuss the implication of different choices of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. When choosing $\boldsymbol{\theta}_{i}$ near the boundary of $\Theta_{i-1}$, as shown in Fig. 5 and Fig. 5, the removal of the hypothesis space depends on the specific direction of cut $\mathbf{bd}\,\mathcal{C}_{i}$. Since the direction of the cut $\mathbf{bd}\,\mathcal{C}_{i}$ will be eventually determined by the direction of user correction according to (17), the removal volume of the hypothesis space can be relatively small, as in Fig. 5, or large, as in Fig. 5, depending on the specific direction of the correction.
As the specific direction of user correction is hard to assume, one can intuitively see that a good choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ will be near the center of $\Theta_{i-1}$. As such, it can lead to an averaged large cut of the hypothesis space regardless of the direction of $\mathbf{bd}\ \mathcal{C}_{i}$, as shown in Fig. 5. We will rigorously discuss the centering choice of $\boldsymbol{\theta}_{i}$ and the guarantee of volume reduction rate of the hypothesis space in the next section.
Figure 5: Illustration of the different choices of $\boldsymbol{\theta}_{i}$. (a) and (b): When $\boldsymbol{\theta}_{i}$ is chosen near the set boundary, the removal volume depends on the specific direction of $\mathbf{bd}\ \mathcal{C}_{i}$. (c): When $\boldsymbol{\theta}_{i}$ is near the center, the cut volume is large on average regardless of the specific user directional correction.
V Main Results and Implementation Details
In this section, we will present the details for how to select $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ at each learning iteration $i$ in our Safe MPC Alignment algorithm. Based on the previous geometric intuition, our choice of $\boldsymbol{\theta}_{i}$ is to achieve a large averaged reduction rate in the hypothesis space volume for all possible human directional corrections. The reduction rate enables us to establish an exponential convergence of the learning process. As a result, we can find an upper bound of total human corrections for successful learning, i.e., $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$, or identify the misspecification of hypothesis space.
@@ V-A The Choice of $\boldsymbol{\theta}_{i}$
As discussed previously, $\boldsymbol{\theta}_{i}$ needs to be picked from the "center" of $\Theta_{i-1}$ in each learning iteration. One option is to choose the mass center of $\Theta_{i-1}$. However, computing the centroid of a polytope can be #P-Hard [53]. While methods like the Hit-And-Run algorithm [54] estimate the mass center by sampling, they can suffer from high sample complexity. Another possible choice is to use the center of the maximum inscribed ball of a polytope [55], but it lacks of the guarantee in volume reduction rate. For computational efficiency and guaranteed convergence, we will choose $\boldsymbol{\theta}_{i}$ as the center of the Maximum Volume Ellipsoid (MVE) inscribed in the convex set $\Theta_{i-1}$, shown in Fig. 6.
@@
=$\mathcal{E}=\{\bar{H}\boldsymbol{\theta}+\bar{\boldsymbol{d}}\ |\ \|\boldsymbol{\theta}\|_{2}\leq 1\},$ (19)
where the affine transformation parameter $\bar{H}\in\mathbb{S}^{r}_{++}$ (a $r\times r$ symmetric positive definite matrix), and $\bar{\boldsymbol{d}}\in\mathbb{R}^{r}$ is called the center of $\mathcal{E}$. $\bar{H},\bar{\boldsymbol{d}}$ are determined via the following optimization problem to maximize the volume of the ellipsoid $\det H$:
$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\max}\quad&\log\det H\\ \rm{s.t.}\quad&\sup\nolimits_{\|\boldsymbol{\theta}\|_{2}\leq 1}\mathbbm{1}_{\Theta}(H\boldsymbol{\theta}+\boldsymbol{d})\leq 0,\end{split}$ (20)
where the indicator function $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=0$ for $\boldsymbol{\theta}\in\Theta$ and otherwise $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=+\infty$. Specifically, when $\Theta$ is a polytope $\{\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{n}_{i}\leq c_{i},\ i=1,\dots,m\}$, the problem can be converted to a convex optimization with $m$ constraints, see chapter 8.4.2 in [56]:
$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\min}\quad&\log\det H^{-1}\\ \rm{s.t.}\quad&\left\lVert H\boldsymbol{n}_{i}\right\rVert_{2}+\boldsymbol{d}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{n}_{i}\leq c_{i},\ i=1,...,m.\end{split}$ (21)
Recall that in our Safe MPC Alignment the learning update (13) is to iteratively cut the hypothesis space by a pair of hyperplanes (17), and that the initial hypothesis space (11) is a box. By induction, it leads to that any updated hypothesis space $\Theta_{i}$ is a convex polytope, $i=0,1,2,...$. Thus, for any $\Theta_{i-1}$, by the above definition, let $\mathcal{E}_{i-1}$ denote the MVE of $\Theta_{i-1}$ and $\bar{\boldsymbol{d}}_{i-1}$ denote the center of $\mathcal{E}_{i-1}$. Then, the choice of $\boldsymbol{\theta}_{i}$ is
@@
The convex optimization (20) can be solved efficiently [57].
@@ V-B Result 1: Finite Iteration Convergence
When choosing $\boldsymbol{\theta}_{i}$ as the center of MVE of $\Theta_{i-1}$, the following lemma states that the volume ratio ${\mathbf{Vol}(\Theta_{i})}/{\mathbf{Vol}(\Theta_{i-1})}$ after each hypothesis space cut can be upper bounded.
@@
In the Safe MPC Alignment algorithm, when choosing $\boldsymbol{\theta}_{i}\in\mathbb{R}^{r}$ as the center of MVE of $\Theta_{i-1}$, one has
$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r},$ (23)
@@
See Appendix A-D. ∎
The lemma above asserts an upper bound of the volume of the updated hypothesis space $\mathbf{Vol}(\Theta_{i})$ relative to the previous volume. With this, we can now establish the finite iteration convergence of the Safe MPC Alignment, stated below.
@@
In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center of MVE of $\Theta_{i-1}$, $i=1,2,3,...$. If $\bar{\Theta}_{H}\subset\Theta_{0}$, the algorithm will terminate successfully within
@@
See Appendix A-E. ∎
The above main result asserts a maximum iteration count for the proposed algorithm to find $\boldsymbol{\theta}_{i}\in\Theta_{H}$. In other words, within $K$ rounds of human corrections, the robot can successfully learn a safety constraint intended by the user. The above result requires a pre-set of $\rho_{H}^{r}$, which is the radius of a ball contained in $\bar{\Theta}_{H}$. Since the size of $\bar{\Theta}_{H}$ is unknown in practice, $\rho_{H}>0$ can be interpreted as an underestimated size of the unknown user intent set $\bar{\Theta}_{H}$. It also serves as a termination threshold for learning convergence: a smaller $\rho_{H}$ can lead to a more concentrated of $\boldsymbol{\theta}_{H}$ learned at convergence.
However, the finite iteration convergence in Theorem 1 comes with a prerequisite: $\bar{\Theta}_{H}\in\Theta_{0}$ has to be satisfied in the first place in the Safe MPC Alignment algorithm. Although such a condition can be achieved by empirically choosing $\Theta_{0}$ large enough to contain all possible $\bar{\Theta}_{H}$ (10), it is also desired to have a guaranteed method to detect the hypothesis misspecification, where such condition is not met. Next, we will discuss the declaration of hypothesis misspecification.
@@ V-C Result 2: Hypothesis Misspecification Declaration
We consider the case of hypothesis space misspecification, where $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$. In this case, no $\boldsymbol{\theta}\in\Theta_{0}$ can account for the user intended safety (4). As shown in [52, 58], the hypothesis misspecification can lead to the undesired behavior of the robot. The following result states the proposed Safe MPC Alignment algorithm can automatically certify the misspecification of the initial hypothesis space.
Theorem 2 (Certifying misspecification).
In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center of MVE of $\Theta_{i-1}$, $i=1,2,3,...$. If $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the algorithm converges to a set $\Theta^{*}$ as $i\rightarrow\infty$, such that $\mathbf{Vol}(\Theta^{*})=0$ and $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$.
@@
See Appendix A-F. ∎
The above result states that when the initial hypothesis space is misspecified, $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the Safe MPC Alignment algorithm will finally converge to a set of zero volume on the boundary of the initial hypothesis space $\Theta_{0}$.
The above theorem inspires us to empirically identify the misspecification by monitoring the position of the MVE center of $\Theta_{i}$ during the learning update. As shown in Fig. 7, if after the maximum iteration $K-1$ in (24), the MVE center $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of the initial hypothesis set $\Theta_{0}$, there is likely hypothesis space misspecification, and the algorithm terminates with misspecification failure.
Figure 7: When there is misspecification, the Safe MPC Alignment algorithm will converge to some point on the boundary of $\Theta_{0}$. So we can use the distance of $\boldsymbol{\theta}_{K-1}$ to the boundary as an indicator for misspecification.
Specifically, recall that $\Theta_{0}$ in (10) is a box with bounds $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$, we can pre-define a small threshold $\epsilon$ to examine if $\boldsymbol{\theta}_{K-1}$ is on the boundary of $\Theta_{0}$. If $\boldsymbol{\theta}_{K-1}$ satisfies
$\min_{j{=}1,...,r}\ \min\big{(}\lvert\boldsymbol{c}_{l}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert,\lvert\boldsymbol{c}_{u}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert\big{)}\leq\epsilon$ (25)